# PsOCR: Benchmarking Large Multimodal Models for Optical Character Recognition in Low-resource Pashto Language

### v1

@@ Abstract
This paper evaluates the performance of Large Multimodal Models (LMMs) on Optical Character Recognition (OCR) in the low-resource Pashto language. Natural Language Processing (NLP) in Pashto faces several challenges due to the cursive nature of its script and a scarcity of structured datasets. To address this, we developed a synthetic Pashto OCR dataset, PsOCR, consisting of one million images annotated with bounding boxes at word, line, and document levels, suitable for training and evaluating models based on different architectures, including Convolutional Neural Networks (CNNs) and Transformers. PsOCR covers variations across 1,000 unique font families, colors, image sizes, and layouts. A benchmark subset of 10K images was selected to evaluate the performance of several LMMs, including seven open-source models: DeepSeek’s Janus, InternVL, MiniCPM, Florence, and Qwen (3B and 7B), and four closed-source models: GPT-4o, Gemini, Claude, and Grok. Experimental results demonstrate that Gemini achieves the best performance among all models, whereas among open-source models, Qwen-7B stands out. This work provides an insightful assessment of the capabilities and limitations of current LMMs for OCR tasks in Pashto and establishes a foundation for further research not only in Pashto OCR but also for other similar scripts such as Arabic, Persian, and Urdu. PsOCR is available at https://github.com/zirak-ai/PashtoOCR.
Keywords OCR  $\cdot$ LMMs  $\cdot$ VLMs  $\cdot$ NLP  $\cdot$ Datasets  $\cdot$ Benchmarks  $\cdot$ Computer Vision
Figure 1: Performance Comparison of the top-6 LMMs on Pashto OCR
@@ 1 Introduction
Synthetic data has emerged as an invaluable resource for addressing data scarcity issues in OCR research, particularly for low-resource languages and scripts. Generating synthetic datasets allows researchers to develop and benchmark robust AI models without the substantial time and resource investments required for manual annotation. This approach has proven effective across diverse languages and scripts; for instance, [14] demonstrated that incorporating synthetic data alongside real-world data can lead to performance improvements. Similarly, [15] introduced a comprehensive synthetic dataset for 23 Indic languages, significantly improving accuracy through fine-tuning on synthetic samples. Moreover, [16] applied synthetic data generation techniques for Arabic OCR, facilitating extensive system evaluation and comparative studies. These studies underscore the importance and effectiveness of synthetic data in enhancing OCR model performance.
In this work, we introduce PsOCR, a comprehensive dataset for training and evaluating LMMs on Pashto OCR tasks. PsOCR is composed of one million synthetic images, annotated at word, line, and document levels, featuring around a thousand unique font families, varied color schemes, and diverse layouts. A curated subset of 10K images is used as an evaluation benchmark to examine the performance of various flagship LMMs. We have evaluated a total of eleven models, including seven open-source models: InternVL [17], MiniCPM [18], Llama [19], Janus [20], Florence [21], and Qwen (3B and 7B) [22], and four proprietary models: Grok [23], Claude [24], GPT-4o [25], and Gemini [26]. Detailed experimental analyses reveal each model’s strengths and weaknesses in Pashto text extraction, paving the way for further research in this domain.
@@ 2.1 Pashto OCR
One of the foundational works in Pashto OCR is [10], identifying the primary recognizable units in Pashto script. They suggested “ligature” as the basic unit for OCR and claimed that 7,681 basic shapes were adequate for representing all Pashto ligatures, thus simplifying recognition challenges. [27] proposed a robust OCR method addressing scale, rotation, and location invariances in Pashto script; they developed a holistic recognition framework and created a dataset of 8K images covering 1K unique ligatures. Recognizing dataset availability as a significant bottleneck, [28] introduced the KPTI database containing 17K handwritten and printed text-line images, facilitating benchmarking of deep-learning models. [29] used Sequential Minimal Optimization with horizontal and vertical projections for line, word, and character segmentation and Local Binary Patterns for feature extraction; this model achieved satisfactory performance and emphasized the importance of precise segmentation and robust feature extraction. [30] leveraged transfer learning with pre-trained CNNs, highlighting the efficacy of fine-tuned deep-learning models augmented by data-enhancement techniques. [31] developed a CNN-based classifier specifically targeting handwritten numerals, capable of handling variations in style and orientation. Similarly, [32] and [33] used CNNs and developed their own benchmark datasets for evaluation. [34] tested feed-forward neural networks with varying ReLU layers on a custom dataset, providing crucial insights into neural architectures, while [32] used LSTM for character recognition. [35] developed the publicly accessible HPCID dataset, comprising 15K handwritten samples, offering critical resources for training and evaluation. Lastly, [36] created the PHTI dataset, encompassing 36K segmented text-line images from diverse genres and writers, filling a substantial gap in available Pashto handwritten resources.
@@ 2.3 OCR Benchmarks for LMMs
Under the large umbrella of OCR, various benchmarks are available for different tasks, such as for Visual Question Answering (VQA): TextOCR [41], TextCaps [42], ST-VQA [43], OCR-VQA [44], TextVQA [45], DocVQA [46], InfographicVQA [47], ChartQA [48], MTVQA [49], MultipanelVQA [50], EST-VQA [51]; for rich-text image understanding: LLaVAR [52], MMR [53]; and for reasoning: MM-GNN [54], textKVQA [55], OCRBench [56]. Various other benchmarks are available for specific purposes, such as KITAB-Bench [57] and CAMEL-Bench [58] for Arabic OCR tasks, MOTBench [59] for menu understanding and translation, CC-OCR [60] for literacy, and Fox [61] for fine-grained and multi-page document understanding. However, there is no significant prior work comparable to ours, which specifically focuses on benchmarking LMMs for text extraction in the low-resource Pashto language.
@@ 3.1 Text Corpus Collection
Pashto has limited structured textual content available for large-scale NLP and OCR tasks. Developing an OCR dataset at a significant scale requires a substantial and high-quality raw text corpus. To address this requirement, we collected Pashto text from three primary sources. Firstly, we attempted extraction from the Common Crawl (CC)  [^1] Common Crawl: https://commoncrawl.org corpus. Although CC is extensive, the availability of Pashto text is extremely low; filtering a 15,000 GB dataset specifically for Pashto (ISO-639-3 language code: “pus”) yielded only $\approx$1GB of text, constituting about 0.008% of the entire corpus. This indicates that relying solely on the CC corpus is inefficient for large-scale Pashto text collection. Secondly, we crawled open-source websites abundant in Pashto content, significantly supplementing our corpus. Third, we incorporated existing text resources from Twitter, books, and news websites used in previous studies [11], [62], [63]. By combining these three data sources, we created a text corpus, sufficient for building a comprehensive OCR dataset.
@@ 4.1 Models Selection
In total, this study evaluates eleven LMMs: seven open-source (InternVL, MiniCPM, Llama, Janus, Florence, Qwen 3B, and Qwen 7B) and four proprietary (GPT-4o, Gemini, Claude, and Grok). A concise overview of these models, along with their parent organizations, specific variants, and parameter counts, is summarized in Table 1.
@@
InternVL [17] OpenGVLab InternVL2_5-8B 8.08B
MiniCPM [18] OpenBMB MiniCPM-o-2_6 8.67B
@@
Grok [23] X-AI grok-2-vision-1212 -
Claude [24] Anthropic claude-3-7-sonnet-20250219 -
GPT-4o [25] OpenAI GPT-4o (2024-08-06) -
Gemini [26] Google gemini-2.0-flash -
Table 1: Details of the Evaluated Models
@@ 4.2 Experimental Setup
The core objective of this study is to assess the zero-shot OCR performance of selected LMMs on our newly developed PsOCR benchmark. Neither pre-training nor fine-tuning was applied; models were evaluated directly in their original form. The inference pipeline for proprietary models: GPT-4o, Gemini, Claude, and Grok was developed through their respective APIs. Conversely, selected open-source models, including InternVL, MiniCPM, Janus, Llama, and Qwen, were downloaded and tested locally. Decisions regarding the use of APIs versus local inference were driven by model accessibility and available computational resources. For the experiment, we used Python-based tools and libraries such as openai [^5] OpenAI Python API library: https://github.com/openai/openai-python, HuggingFace’s Transformers [^6] HuggingFace Transformers: https://github.com/huggingface/transformers, Google’s GenAI SDK [^7] Google Gen AI SDK: https://github.com/googleapis/python-genai, and Anthropic’s API [^8] Anthropic API: https://docs.anthropic.com/en/api/getting-started. Local inference ran on a PC with an Intel Core i7 processor, 32GB of RAM, and an NVIDIA RTX 4080 GPU. During evaluation, each of the 10K images was provided one by one to each model along with a specifically engineered prompt, and the outputs were recorded. Although most evaluated models support conversational context, our experiment explicitly avoided using histories; the evaluation strictly adhered to a zero-shot setting, providing no contextual examples or prior information, to ensure an unbiased assessment of their innate OCR capabilities.
@@ 4.3 Prompt Engineering
A critical factor affecting evaluation results in zero-shot scenarios is the quality and specificity of the instructions provided to the models, known as “prompt.” Effective prompt design is essential to achieving accurate model responses. However, no universal standards or guidelines exist for crafting optimal prompts. Different models have been trained and fine-tuned by diverse research teams using varying methodologies, leading to notable differences in how models interpret and respond to prompts. Consequently, a prompt effective for one model may not be equally suitable for another. To address this, we invested considerable effort in designing customized prompts for each model, adopting a trial-and-error approach to iteratively refine instructions and maximize performance. A primary difficulty encountered during prompt engineering involved clearly instructing the models to return only the exact text displayed within the images, without providing additional contextual information, explanations, or translations. Several models demonstrated a tendency to translate the extracted text or include extraneous details. Thus, significant prompt refinement was necessary to explicitly discourage translation and ensure that the outputs precisely reflected the original Pashto text. Figure 7 and Figure 8 present two prompt examples that yielded the best results in our experiments.
@@
Figure 8: Prompt for all the other models, including GPT4-o, Gemini, and Qwen
@@ 4.4 Post Processing Models’ Responses
4.5 Evaluation Matrices
To quantitatively assess the OCR capabilities of the evaluated LMMs, we employed three well-established performance metrics: Character Accuracy, Word Accuracy, and BLEU score, and also computed the average (Avg) of these three metrics.
4.5.1 Character Accuracy
Character Accuracy (CA) measures the accuracy at the character level and is derived from the Character Error Rate (CER), as follows:
$CA=1-CER$ (1)
CER itself is calculated by counting the minimum number of character-level operations (insertions, deletions, substitutions) required to transform the predicted text string into the corresponding reference text string (ground-truth text), normalized by the total number of characters in the reference text:
CER=Number of Character Errors (Insertions + Deletions + Substitutions)Total Number of Characters in Reference $\mbox{CER}=\frac{\mbox{Number of Character Errors (Insertions + Deletions + Substitutions)}}{\mbox{Total Number of Characters in Reference Text}}$
4.5.2 Word Accuracy
Word Accuracy (WA) evaluates accuracy at the word level and is derived similarly from the Word Error Rate (WER), as follows:
$WA=1-WER$ (2)
The WER calculation involves counting the minimum number of word-level operations (insertions, deletions, substitutions) required to convert the predicted text into the corresponding reference text, normalized by the total number of words in the reference text:
WER=Number of Word Errors (Insertions + Deletions + Substitutions)Total Number of Words in Reference $\mbox{WER}=\frac{\mbox{Number of Word Errors (Insertions + Deletions + Substitutions)}}{\mbox{Total Number of Words in Reference Text}}$
4.5.3 BLEU Score
The BLEU score measures the similarity of the predicted text compared to the reference text based on matching word n-grams. A higher BLEU score indicates greater similarity to the reference text; a detailed explanation can be found in [64].
4.5.4 Average
To compress each model’s overall performance into a single scalar for easier comparison, we compute an Average (Avg) score by equally weighting CA, WA, and BLEU for each test image, and then averaging these per-image scores across the entire benchmark. Formally:
$Avg=\frac{1}{n}\sum_{i=1}^{n}\frac{(CA_{i}+WA_{i}+BLEU_{i})}{3}$ (3)
Where $n$ is the total number of images in the benchmark (10K).

### v2

@@ Abstract
This paper evaluates the performance of Large Multimodal Models (LMMs) on Optical Character Recognition (OCR) for the low-resource Pashto language. Pashto OCR is challenging due to its cursive Perso-Arabic script and the scarcity of large-scale annotated datasets. To address these challenges, we introduce PsOCR, a large-scale synthetic Pashto OCR dataset containing one million images annotated at the word, line, and document levels. PsOCR includes extensive variability across 1,000 font families, font sizes, colors, image resolutions, and layouts. A benchmark subset of 10,000 images is used to evaluate several state-of-the-art LMMs, including Llama, Florence, Qwen-3B/7B, GPT-4o, Gemini, Claude, and Grok, under zero-shot settings. Experimental results demonstrate that Gemini achieves the best overall performance, while Qwen-7B stands out among open-source models. This work provides valuable insights into the capabilities and limitations of current LMMs for Pashto OCR and establishes a foundation for future research in languages with similar scripts.
keywords:
OCR, Benchmarks, Computer Vision, Datasets, Image Processing, LLMs, NLP, Pashto, Vision-Language Models
[…]
@@ 1 Introduction
Synthetic data has emerged as an invaluable resource for addressing data scarcity issues in OCR research, particularly for low-resource languages and scripts. Generating synthetic datasets allows researchers to develop and benchmark robust AI models without the substantial time and resource investments required for manual annotation. This approach has proven effective across diverse languages and scripts; for instance, [14] demonstrated that incorporating synthetic data alongside real-world data can lead to performance improvements. Similarly, [15] introduced a comprehensive synthetic dataset for 23 Indic languages, significantly improving accuracy through fine-tuning on synthetic samples. Moreover, [16] applied synthetic data generation techniques for Arabic OCR, facilitating extensive system evaluation and comparative studies. These studies underscore the importance and effectiveness of synthetic data in enhancing OCR model performance. In this work, we introduce PsOCR, a comprehensive dataset for training and evaluating LMMs on Pashto OCR tasks. PsOCR is composed of one million synthetic images, annotated at word, line, and document levels, featuring around a thousand unique font families, varied color schemes, and diverse layouts. A curated subset of 10K images is used as an evaluation benchmark to examine the performance of various flagship LMMs. We have evaluated a total of seven models, including four open-source models: Llama [19], Florence [21], and Qwen (3B and 7B) [22], and four proprietary models: Grok [23], Claude [24], GPT-4o [25], and Gemini [26]. Detailed experimental analyses reveal each model’s strengths and weaknesses in Pashto text extraction, paving the way for further research in this domain.
@@ 2.1 Pashto OCR
One of the foundational works in Pashto OCR is [10], identifying the primary recognizable units in Pashto script. They suggested “ligature” as the basic unit for OCR and claimed that 7,681 basic shapes were adequate for representing all Pashto ligatures, thus simplifying recognition challenges. [27] proposed a robust OCR method addressing scale, rotation, and location invariances in Pashto script; they developed a holistic recognition framework and created a dataset of 8K images covering 1K unique ligatures. Recognizing dataset availability as a significant bottleneck, [28] introduced the KPTI database containing 17K handwritten and printed text-line images, facilitating benchmarking of deep-learning models. [29] used Sequential Minimal Optimization with horizontal and vertical projections for line, word, and character segmentation and Local Binary Patterns for feature extraction; this model achieved satisfactory performance and emphasized the importance of precise segmentation and robust feature extraction. [30] leveraged transfer learning with pre-trained CNNs, highlighting the efficacy of fine-tuned deep-learning models augmented by data-enhancement techniques. [31] developed a CNN-based classifier specifically targeting handwritten numerals, capable of handling variations in style and orientation. Similarly, [32] and [33] used CNNs and developed their own benchmark datasets for evaluation. [34] tested feed-forward neural networks with varying ReLU layers on a custom dataset, providing crucial insights into neural architectures, while [32] used LSTM for character recognition. [35] developed the publicly accessible HPCID dataset, comprising 15K handwritten samples, offering critical resources for training and evaluation. Lastly, [36] created the PHTI dataset, encompassing 36K segmented text-line images from diverse genres and writers, filling a substantial gap in available Pashto handwritten resources. A summary of isolated character datasets for Pashto OCR is given in Table 1, while Table 2 compares our dataset with previous related works.
Table 1: Summary of isolated characters datasets for Pashto OCR.
Ref. Type Size Classes Granularity Public
Ahmad et al. [10] Synthetic 8K 1,000 ligatures Ligature Yes
FAST-NU [30] Synthetic 4K 1,000 ligatures Ligature Yes
Ullah et al. [29] Print 5K - Character No
Uddin et al. [34] Handwritten 43K 43 characters Character Yes
Poha [31] Handwritten 26.4K 44 chars+10 digits Character Yes
Khan et al. [32] Handwritten 4.5K 44 characters Character No
Rehman et al. [33] Handwritten 106K 53 symbols Character Partial
Khan et al. [67] Handwritten 50K 10 digits Digit Yes
Table 2: Comparison of text-level Pashto OCR datasets.
   Attribute    KPTI [28]    PHTI [36]    PsOCR (Ours)
   Dataset Type    Real (Scanned)    Real (Handwritten)    Synthetic
   Scale    17K lines    36K lines    5.89M Lines/1M images
   Font Color Variation    $\times$    $\times$    $\checkmark$
   Bg Color Variation    $\times$    $\times$    $\checkmark$
   Font Family Variation    $\times$    $\times$    $\checkmark$
   Font Size Variation    $\times$    $\times$    $\checkmark$
   Word-level Annotation    $\times$    $\times$    $\checkmark$
   Line-level Annotation    $\checkmark$    $\checkmark$    $\checkmark$
   Doc-level Annotation    $\times$    $\times$    $\checkmark$
   Writer Diversity    $\times$    $\checkmark$    $\times$
   Skew/Rotation    $\checkmark$    $\checkmark$    $\times$
@@ 2.3 OCR Benchmarks for LMMs
Under the large umbrella of OCR, various benchmarks are available for different tasks, such as for Visual Question Answering (VQA): TextOCR [41], TextCaps [42], ST-VQA [43, 17], OCR-VQA [44], TextVQA [45], DocVQA [46], InfographicVQA [47], ChartQA [48], MTVQA [49], MultipanelVQA [50], EST-VQA [51]; for rich-text image understanding: LLaVAR [52], MMR [53]; and for reasoning: MM-GNN [54, 18], textKVQA [55], OCRBench [56]. Various other benchmarks are available for specific purposes, such as KITAB-Bench [57] and CAMEL-Bench [58] for Arabic OCR tasks, MOTBench [59] for menu understanding and translation, CC-OCR [60] for literacy, and Fox [61, 20] for fine-grained and multi-page document understanding. However, there is no significant prior work comparable to ours, which specifically focuses on benchmarking LMMs for text extraction in the low-resource Pashto language.
@@ 3.1 Text Corpus Collection
Pashto has limited structured textual content available for large-scale NLP and OCR tasks. Developing an OCR dataset at a significant scale requires a substantial and high-quality raw text corpus. To address this requirement, we collected Pashto text from three primary sources. Firstly, we attempted extraction from the Common Crawl (CC)  [^1] http://www.ieee.org/authortools corpus. Although CC is extensive, the availability of Pashto text is extremely low; filtering a 15,000 GB dataset specifically for Pashto (ISO-639-3 language code: “pus”) yielded only $\approx$1GB of text, constituting about 0.008% of the entire corpus. This indicates that relying solely on the CC corpus is inefficient for large-scale Pashto text collection. Secondly, we crawled open-source websites abundant in Pashto content, significantly supplementing our corpus. Third, we incorporated existing text resources from Twitter, books, and news websites used in previous studies [11], [62], [63], which constitute the major portion of our corpus. By combining these three data sources, we created a text corpus, sufficient for building a comprehensive OCR dataset.
@@ 4.1 Models Selection
In total, this study evaluates seven VLMs: four open-source (Llama, Florence, Qwen 3B, and Qwen 7B) and four proprietary (GPT-4o, Gemini, Claude, and Grok). A concise overview of these models, along with their parent organizations, specific variants, and parameter counts, is summarized in Table 3.
Table 3: Summary of the Evaluated Models
@@
  Grok [23]   X-AI   grok-2-vision-1212   –
  Claude [24]   Anthropic   claude-3-7-sonnet-20250219   –
  GPT-4o [25]   OpenAI   GPT-4o (2024-08-06)   –
  Gemini [26]   Google   gemini-2.0-flash   –
@@ 4.2 Experimental Setup
The core objective of this study is to assess the zero-shot OCR performance of selected LMMs on our newly developed PsOCR benchmark. Neither pre-training nor fine-tuning was applied; models were evaluated directly in their original form. The inference pipeline for proprietary models: GPT-4o, Gemini, Claude, and Grok was developed through their respective APIs. Conversely, selected open-source models, including Llama, Florence, and Qwen, were downloaded and tested locally. Decisions regarding the use of APIs versus local inference were driven by model accessibility and available computational resources. For the experiment, we used Python-based tools and libraries such as openai [^5] OpenAI Python API library: https://github.com/openai/openai-python, HuggingFace’s Transformers [^6] HuggingFace Transformers: https://github.com/huggingface/transformers, Google’s GenAI SDK [^7] Google Gen AI SDK: https://github.com/googleapis/python-genai , and Anthropic’s API [^8] Anthropic API: https://docs.anthropic.com/en/api/getting-started. Local inference ran on a PC with an Intel Core i7 processor, 32GB of RAM, and an NVIDIA RTX 4080 GPU. During evaluation, each of the 10K images was provided one by one to each model along with a specifically engineered prompt, and the outputs were recorded. Although most evaluated models support conversational context, our experiment explicitly avoided using histories; the evaluation strictly adhered to a zero-shot setting, providing no contextual examples or prior information, to ensure an unbiased assessment of their innate OCR capabilities.
@@ 4.3 Prompt Engineering
A critical factor affecting evaluation results in zero-shot scenarios is the quality and specificity of the instructions provided to the models, known as “prompt.” Effective prompt design is essential to achieving accurate model responses. However, no universal standards or guidelines exist for crafting optimal prompts. Different models have been trained and fine-tuned by diverse research teams using varying methodologies, leading to notable differences in how models interpret and respond to prompts. Consequently, a prompt effective for one model may not be equally suitable for another. To address this, we invested considerable effort in designing customized prompts for each model. First we wrote several prompts following some standard guidlines, such as those mentioned in [64, 65]. Then we chose the prompt giving the best performance. After selecting the best prompts we adopted a trial-and-error approach to iteratively refine instructions to further maximize performance. A primary difficulty encountered during prompt engineering involved clearly instructing the models to return only the exact text displayed within the images, without providing additional contextual information, explanations, or translations. Several models demonstrated a tendency to translate the extracted text or include extraneous details. Thus, significant prompt refinement was necessary to explicitly discourage translation and ensure that the outputs precisely reflected the original Pashto text. Figure 6 and Figure 7 present two prompt examples that yielded the best results in our experiments.
@@
Figure 7: Prompt for GPT4-o, Gemini, Grok, Florence and Qwen
@@ 4.4 Post Processing Models’ Responses
4.5 Evaluation Metrics
To comprehensively evaluate the document-level OCR performance of LMMs, we adopt two complementary categories of evaluation metrics: (1) transcription accuracy metrics and (2) text similarity metrics. This dual evaluation strategy allows us to assess both the exactness of OCR transcription and the high-level similarity between ground-truth and predicted texts.
4.5.1 Transcription Accuracy Metrics
Transcription accuracy metrics are standard in OCR research and measure how precisely the recognized text matches the ground-truth transcription at the character and word levels. These metrics are based on edit distance and explicitly penalize substitution, insertion, and deletion errors, making them well-suited for evaluating OCR performance in low-resource and cursive scripts such as Pashto. We employ the following two widely used transcription accuracy metrics.
Character Error Rate (CER)
Character Error Rate measures the normalized edit distance between the predicted text and the ground-truth text at the character level. It is defined as:
$\mathrm{CER}=\frac{S+D+I}{N},$ (1)
where $S$, $D$, and $I$ denote the number of character substitutions, deletions, and insertions, respectively, and $N$ is the total number of characters in the ground-truth text. Lower CER values indicate better OCR performance.
Word Error Rate (WER)
Word Error Rate evaluates transcription errors at the word level and is computed analogously to CER:
$\mathrm{WER}=\frac{S_{w}+D_{w}+I_{w}}{N_{w}},$ (2)
where $S_{w}$, $D_{w}$, and $I_{w}$ represent word-level substitutions, deletions, and insertions, and $N_{w}$ is the total number of words in the ground-truth document. WER reflects word-level usability of OCR output for downstream NLP applications.
4.5.2 Text Similarity Metrics
In addition to transcription accuracy metrics, we report text similarity metrics to analyze the overall similarity between predicted and ground-truth documents. Since this study evaluates transformer-based generative models, such metrics provide complementary insights into the models’ ability to generate text sequences that are globally similar to the reference, even when minor transcription errors are present. These metrics are reported as auxiliary measures and are not treated as direct indicators of OCR accuracy.
BLEU
BLEU measures the precision of overlapping $n$-grams between the predicted text and the ground-truth text. It is computed as:
$\mathrm{BLEU}=\mathrm{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (3)
where $p_{n}$ denotes the modified $n$-gram precision, $w_{n}$ are weighting factors, and $\mathrm{BP}$ is a brevity penalty that penalizes overly short predictions.
METEOR
METEOR evaluates text similarity based on unigram alignment between the predicted and reference texts, incorporating both precision and recall. It is defined as:
$\mathrm{METEOR}=F_{\text{mean}}\cdot(1-P),$ (4)
where $F_{\text{mean}}$ is the harmonic mean of unigram precision and recall, and $P$ is a penalty term that accounts for fragmented alignments.
Bag-of-Words (BoW) Similarity
BoW evaluation measures text similarity by comparing the unordered sets of words in the predicted and ground-truth texts, ignoring word order. Each document is represented as a word frequency vector, and similarity is computed using cosine similarity:
$\mathrm{BoW}=\frac{\mathbf{v}_{\text{gt}}\cdot\mathbf{v}_{\text{pred}}}{\|\mathbf{v}_{\text{gt}}\|\,\|\mathbf{v}_{\text{pred}}\|},$ (5)
where $\mathbf{v}_{\text{gt}}$ and $\mathbf{v}_{\text{pred}}$ denote the BoW vectors of the ground-truth and predicted texts, respectively.